OpenStudy (anonymous):

can some one show me how to multiply and simplify: (5+2 radical 2)(8-3 radical2)

OpenStudy (anonymous):

\[(5+2\sqrt{2})(8-3\sqrt{?}\]

OpenStudy (anonymous):

FOIL it

OpenStudy (anonymous):

40 - 15√2 +16√2 -12 = 28 + √2

OpenStudy (anonymous):

Alright that made sence haha thank you=)